OpenStudy (anonymous):

9x+8y=-5 -8x+y=45 solve by substitution method

OpenStudy (anonymous):

X= Y=

OpenStudy (mimi_x3):

y = 8x + 45 9x+ 8(8x+45) = -5 9x + 8(8x + 45) = -5 9x + 64x + 360 = -5 73x = - 365 x = -5 Substitute x =-5 into 9x + 8y = - 5 -45 + 8y = - 5 8y = 40 y = 5

OpenStudy (anonymous):

or from 9x+8y=-5 8y=-9x-5 y=(-9x-5)/8 sub to -8x+y=45 to get -8x+(-9x-5)/8 =45 , mult by 8 on both sides to get -64x+(-9x-5)=360 -73x-5=360 -73x=360+5=365 x=-365/73=-5 ans, sub this to 8y=-9x-5 8y=-9(-5)-5 8y=45-5=40 y=40/8=5 ans

OpenStudy (anonymous):

my pc keeps hanging up on me..lol

OpenStudy (anonymous):

do you wanna know how you check if you work is correct?

OpenStudy (anonymous):

there is another method that you need to learn, you can use this also for checking your work..by addition or subtraction simultaneously

OpenStudy (anonymous):

9x+8y=-5 9x+8y=-5 -8x+y=45 mult by 8,then subtract -( -64x+8y=360) ------------- 73x +0=-365 x=-365/73=-5 now sub this to the other EQ

OpenStudy (anonymous):

good luck now stacy,,have fun..lol

OpenStudy (anonymous):

thanks